### v5

# The AI Productivity Index: APEX-v1-extended

Abstract
We present an extended version of the AI Productivity Index (APEX-v1-extended), a benchmark for assessing whether frontier models are capable of performing economically valuable tasks in four jobs: investment banking associate, management consultant, big law associate, and primary care physician (MD). This technical report details the extensions to APEX-v1, including an increase in the held-out evaluation set from $n=50$ to $n=100$ cases per job ($n=400$ total) and updates to the grading methodology. We present a new leaderboard, where GPT5 (Thinking = High) remains the top performing model with a score of $67.0\%$. APEX-v1-extended shows that frontier models still have substantial limitations when performing typical professional tasks. To support further research, we are open sourcing $n=25$ non-benchmark example cases per job ($n=100$ total) along with our evaluation harness.
Figure 1: Models’ mean score on APEX-v1-extended. The labels in parentheses show the “Thinking” setting.
Figure 2: Production process for creating APEX-v1-extended.
1 Introduction
AI benchmarks give insight into models’ strengths and weaknesses, helping to gauge their capabilities and compare performance. Good benchmarking is essential for effective hill climbing and progress in AI (Kiela et al., 2021; Zhang et al., 2024; Schwartz et al., 2025; Weidinger et al., 2025). However, many benchmarks are scholastic, abstract and narrow, and uncoupled from real-world applications. With a few exceptions (such as GDPVal from OpenAI (OpenAI, 2025) and BigLaw Bench from Harvey (Team, 2024)), most benchmarks do not test frontier AI systems for their ability to deliver real world value, such as improving the productivity of human workers (Cahn, 2024; Bar-Gill and Sunstein, 2025; Filippucci et al., 2024; Tamkin and McCrory, 2025). In September 2025 we introduced the AI Productivity Index (APEX-v1.0) to tackle this problem Vidgen et al. (2025), bridging the gap between what professionals want from AI systems and what benchmarks test for. We now introduce APEX-v1-extended. It expands the breadth and robustness of APEX and has an improved grading methodology.
APEX-v1-extended contains realistic and varied prompts that reflect the everyday work of professionals. They are split across the same four knowledge jobs as the original v1 benchmark: investment banking associate, management consultant, big law associate, and primary care physician (MD). We have increased the size of the hidden heldout set from $n=50$ to $n=100$ per job, increasing the overall size of the benchmark from $n=200$ to $n=400$. The new cases have the same design, comprising a prompt, source documents that the model has to reason over, and a prompt-specific grading rubric. See an example in Figure  3. The $n=400$ cases will remain a closed heldout dataset for rigorous evaluation of frontier models. We are also open sourcing a dataset of $n=25$ cases per job ($n=100$ total) with a $C$-BY license. We refer to this as apex-v1-devset. [^1] APEX-v1-extended on Hugging Face. Our eval harness is available on Github. [^2] Mercor on Github.
Changes to grading in APEX-v1-extended include: (1) switching the panel of LM judges to a single LM, Gemini 2.5 Flash, (2) passing the prompts to each model 8 times and taking a mean (rather than 3 times and taking a median) and (3) reporting 95% confidence intervals with model scores. We have used this methodology to create a new leaderboard for APEX-v1-extended with results for 10 frontier models.333mercor.com/apex. GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. APEX-v1-extended shows that models struggle with real-world tasks that professionals undertake everyday.
Table 1: Overview of the APEX-v1-extended dataset. For the heldout evalset ($n=100$ per job) and the open source data ($n=25$ per job) we show the mean number of criteria and sources per case and the mean number of tokens in the prompts. Note that the heldout eval set is a recalibrated version of the original set of $n=50$ with $n=50$ new cases.
Dataset Job
No. of
cases
No. of
criteria
No. of
sources
No. of tokens
in prompt
Heldout set Medicine 100 18.39 3.77 177
Open source Medicine 25 16.76 1.08 152
Heldout set Law 100 13.98 4.79 407
Open source Law 25 9.88 3.32 367
Heldout set Management consulting 100 14.71 3.21 409
Open source Management consulting 25 9.68 1.16 269
Heldout set Investment banking 100 12.17 3.01 439
Open source Investment banking 25 9.28 1.60 500
Heldout set Mean for all four jobs 100 14.81 3.70 358
Open source Mean for all four jobs 25 11.40 1.79 322
Figure 3: Example prompt and rubric from the Investment banking associate job in APEX-v1-devset, with $10$ criteria. The prompt is passed to models with one source, a financial model in an XLSX file.
2 Dataset overview
The composition of APEX-v1-extended is shown in Table 1 and the production process is shown in Figure  2.
2.1 Selection of experts
From the Mercor platform we sourced experts with extensive professional experience, and prioritized experts with data labeling experience. To join the project, experts completed a 30- to 45-minute interview. Experts who demonstrated excellent domain knowledge and strong communication and reasoning skills were paid to complete a 1-2 hour assessment that tested their ability to write prompts and rubrics. If they successfully completed the assessment, they were contracted to work on APEX-v1.0. Throughout the project we continually checked in with the experts, gave qualitative feedback, and as needed offboarded under-performers. Expert sourcing and vetting did not change from APEX-v1.0 to APEX-v1-extended. The number of experts who worked on the benchmark increased from 76 to 137. Mean industry experience remains at 7+ years.
2.2 Quality control
Experts created each prompt independently, based on their day-to-day work. The mean time for an expert to execute the prompts in a real-world professional setting is estimated at $2.7$ hours, with a range of $0.5$ to 2020 hours. Experts submitted their prompts for review, where they were approved or rejected by reviewers. For rejections, reviewers could either remove the prompt from production or request specific changes. After the prompt was approved, the contributor created the grading rubric, which was also reviewed by a reviewer. Multiple rounds of review help to ensure that cases meet the requirements of the project, are varied in terms of topic, style and model behavior, and are sufficiently challenging. We use in-house LM-powered reviewing tools to give experts immediate feedback on their work. Experts are always responsible for the quality of their submissions. Finally, to ensure that models give a valid single-turn response, we append a short instruction to each prompt, as given in Appendix B.
2.3 Selection of sources
For each case, experts found suitable sources or, if they were not freely available, created them. Sources with restrictive licenses or access controls, such as being behind a paywall, were explicitly prohibited. APEX-v1-extended contains pdf, xlsx, txt, docx and csv files, up to a maximum combined length of 100,000 tokens. [^4] Checked against OpenAI’s tokenizer for GPT 4o. Sources were parsed using Reducto and appended to the prompt in the context window.555reducto.ai The number of sources in APEX-v1-extended is substantially less than in APEX-v1.0, falling from a mean of $5.83$ to $3.70$ in the APEX-v1-extended and $1.79$ in APEX-v1-devset. Experts reported that they could construct realistic prompts by leveraging more content from a single source without the burden of introducing additional sources.
Table 2: Model performance across the four jobs in APEX-v1-extended. Scores represent the mean percentage of criteria met for tasks within each job. GPT 5 (Thinking = High) is the best performing model overall, and the best performing in Law and Medicine. Gemini 3 Pro (Thinking = High) is the best performing model for Investment banking and Management consulting.
Model Provider Overall Invest. banking Law Mgmt. consulting Medicine
Opus 4.5 (On) Anthropic 63.1% 55.2% 74.0% 58.4% 64.6%
Sonnet 4.5 (On) Anthropic 57.2% 45.7% 72.4% 50.7% 59.6%
Opus 4.1 (On) Anthropic 51.4% 42.0% 61.8% 49.4% 52.1%
Gemini 3 Pro (High) Google 64.3% 63.0% 68.5% 64.0% 61.6%
Gemini 2.5 Pro (On) Google 59.4% 54.1% 68.5% 58.9% 55.8%
Gemini 2.5 Flash (On) Google 57.3% 51.1% 68.8% 55.6% 53.6%
GPT 5.1 (High) OpenAI 59.4% 44.3% 77.4% 51.9% 64.0%
GPT 5 (High) OpenAI 67.0% 61.3% 77.9% 63.1% 65.5%
o3 (On) OpenAI 63.5% 57.7% 75.5% 59.0% 61.5%
Grok 4 xAI 63.5% 59.6% 70.2% 59.8% 64.3%
2.4 Rubric creation
To scalably grade each model response, experts created a rubric of quality criteria (Saad-Falcon et al., 2024; Arora et al., 2025; Starace et al., 2025). Rubrics decompose the hard-to-measure concept of response “quality” into testable discrete components. Each criterion is an objective, specific, and self-contained statement about the response, phrased as a descriptive claim. They are analogous to unit tests for code, and can be assessed as Pass or Fail by either a judge LM or a human reviewer. The mean number of criteria per case is $14.81$ in APEX-v1-extended, compared with $11.40$ in APEX-v1-devset. This is a reduction from $29.09$ criteria in APEX-v1.0, due to (1) new cases being added with fewer criteria and (2) excess criteria being removed from the original cases.
3 Leaderboard setup
3.1 Model selection
We tested 10 frontier models from Anthropic, Google DeepMind, OpenAI, and xAI on APEX-v1-extended. Responses were collected at the end of November 2025 from the models’ respective APIs. Thinking is turned On if available and set to High if configurable. If temperature can be configured, we set it to $0.7$. For each prompt, we collect model responses 8 times. This is an increase from 3 times in APEX-v1.0, allowing us to calculate the mean (rather than median) and confidence intervals.
3.2 Grading responses with an LM judge
LM judges are widely used to scalably assess the quality of model responses, supporting high-fidelity evals that can be re-run with no marginal human work (Baumann et al., 2025; Gu et al., 2025; Zhu et al., 2025). LM judging is well-suited for grading rubrics as each criterion is a short self-contained statement. In principle, these statements are easier to judge than overall quality or abstract concepts like “usefulness” or “helpfulness”. We used a single judge in APEX-v1-extended (Gemini 2.5 Flash) rather than the panel of judges for APEX-v1.0. This reduces the eval time, is more transparent, and maintains high performance. The prompt used by the LM judge remains the same, as given in Appendix A.
4 Results
4.1 Leaderboard results
For each model, we compute the mean score from their eight runs on each task. This gives a more reliable measure of model performance. For all 10 models, the mean standard deviation over the eight runs is $9.05$ percentage points, ranging from $8.25$ (Claude Opus 4.5 (Thinking = On)) to $9.5$ (Grok 4).
GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. Models’ mean scores on APEX-v1-extended are shown in Figure 1 with $95\%$ confidence intervals. The differences between models’ scores are statistically significant overall using a Friedman omnibus test ($p<0.000001$). Models’ scores vary across jobs. They are lowest for Investment banking, with the highest scoring model achieving $63.0\%$, followed by Management consulting (top score = $64.0\%$), Medicine (top score = $65.5\%$), and, with substantially higher scores, Law (top score = $77.9\%$). Despite jobs’ differing difficulty, and the different knowledge and reasoning skills they require, models’ rank positions are fairly consistent.
4.2 Performance given task difficulty
Because tasks in APEX-v1-extended are realistic they vary in difficulty, with model scores ranging from $0\%$ to $100\%$. We use paired t-tests on the per-task scores for every pair of models ($n=45$) to directly compare performance. After applying a Bonferroni correction for multiple comparisons, 35 of the 45 pairwise comparisons ($78\%$) remain statistically significant at the adjusted threshold ($p=0.001$). This is also reflected in the z-scores for each model, which indicate how many standard deviations a model’s performance on that task is above or below the mean performance of all models, as shown in Table 3. Z-scores are useful for comparing models but are not suitable for a leaderboard because their values change depending on the set of models that are included. With z-scores, the differences between models are much more stark, showing the substantially better performance of GPT 5 (Thinking = High), which scores 0.50 compared to the second best model, Opus 4.5 (Thinking = On) at 0.28. The rank order of the models changes slightly, with Opus 4.5 (Thinking = On) jumping from 5th to 2nd. This indicates that it is particularly strong on the hardest tasks in APEX-v1-extended.
Table 3: Mean z-scores with 95% confidence intervals for 10 models on APEX-v1-extended.
Model Z-score 95% CI
GPT 5 (High) 0.50 0.42 – 0.58
Opus 4.5 (On) 0.28 0.20 – 0.36
Gemini 3 Pro (High) 0.20 0.10 – 0.29
o3 (On) 0.19 0.11 – 0.27
Grok 4 0.15 0.05 – 0.24
GPT-5.1 (High) -0.00 -0.10 – 0.09
Gemini 2.5 Pro (On) -0.10 -0.18 – -0.02
Sonnet 4.5 (On) -0.18 -0.27 – -0.10
Gemini 2.5 Flash (On) -0.25 -0.34 – -0.16
Opus 4.1 (On) -0.77 -0.86 – -0.68
4.3 Comparison of APEX-v1-extended and APEX-v1-devset
We evaluate the same 10 models against the $n=100$ cases in APEX-v1-devset using the exact same grading methodology (i.e., 8 runs and Gemini 2.5 Flash as a judge). Results are shown in Table 4. GPT 5 (Thinking = High) is the best scoring model on both sets, achieving $65.70\%$ on the APEX-v1-devset versus $67.0\%$ on APEX-v1-extended. Differences in model ranks are minor, with two sets of two models swapped one position, and only one model performs better on APEX-v1-devset than APEX-v1-extended (Gemini 3 Pro (Thinking = High) scoring $1.2\%$ higher). GPT 5.1 (Thinking = High), Gemini 2.5 Flash (Thinking = On) and Sonnet 4.5 (Thinking = On) all score lower on APEX-v1-devset by more than 10 percentage points.
Table 4: Comparison of model scores and ranks between the APEX-v1-extended leaderboard and the APEX-v1-devset.
Model
Bench
OS
Diff
GPT 5 (High) 67.0 (#1) 65.7 (#1) 1.3
Gemini 3 Pro (High) 64.3 (#2) 65.5 (#2) -1.2
Grok 4 63.5 (#3) 61.4 (#3) 2.1
o3 (On) 63.5 (#4) 61.2 (#4) 2.3
Opus 4.5 (On) 63.1 (#5) 54.7 (#5) 8.4
GPT-5.1 (High) 59.4 (#6) 48.2 (#7) 11.3
Gemini 2.5 Pro (On) 59.4 (#7) 53.4 (#6) 6.0
Gemini 2.5 Flash (On) 57.3 (#8) 45.4 (#9) 11.9
Sonnet 4.5 (On) 57.2 (#9) 47.0 (#8) 10.2
Opus 4.1 (On) 51.4 (#10) 43.8 (#10) 7.6
5 Acknowledgments
We are grateful to the expert annotators who contributed to APEX. We thank all our advisors and partners who gave feedback.
Appendix A LM judge
Implementation details:
You are evaluating a model-generated response against a specific criterion. Your task is to determine if the response satisfies this criterion and provide a concise explanation.
Criterion to evaluate: evaluation_case Response to evaluate: model_response
Instructions:
  1. 1.
    First, analyze the response against the given criterion.
  2. 2.
    Determine if the response fully satisfies the criterion (result = 1) or not (result = 0).
  3. 3.
    Provide a concise explanation (maximum 2–3 sentences) that:
    • States whether the criterion is met or not,
    • Points to specific evidence from the response,
    • Avoids unnecessary details or repetition.
Return your evaluation in the following JSON format:
    "result": <1 or 0>,
    "reason": "<concise explanation>"
Keep your explanation brief and focus on the key points that justify your result.
Appendix B Prompt adjustment
Many models are trained to ask followup questions or outline options that the user can confirm or reject. To ensure they give a complete response, we append the following text to every prompt.
**Instruction**
You are an AI assistant that produces final, domain-appropriate deliverables from a given task description and (optionally) attached files. You will be given the following inputs:
  • Task Domain: Domain
  • Task Prompt: Prompt
  • Attachments:
==== Attached files content: ====
  === <File_1> ===
  <File_1_Contents>
  === <File_2> ===
  <File_2_Contents>
  … (repeat as needed)
Ground Rules
  1. 1.
    You must not ask follow-up questions. Interpret the prompt as best you can and produce the best complete answer given the information provided.
  2. 2.
    Use the attachments as primary sources.
  3. 3.
    Treat each "=== <File_Name> === …" block as the full content of that file.
All of the source files that you need have been added to the prompt.